. A driver notices an upcoming speed limit change from 45 mi/h (20 m/s) to 25 mi/h (11 m/s). If she estimates the speed limit will change in 50 m, what acceleration is needed to reach the new speed limit before it begins?


1
Expert's answer
2022-03-08T08:56:22-0500

The acceleration of the car

"a=\frac{v_f^2-v_i^2}{2d}\\
=\rm\frac{(11\:\rm m/s)^2-(20\:\rm m/s)^2}{2*50\: m}=-2.8\:m/s^2"